Identify the concoid 1+2x^2+9y^2=3z^2. Does the XY plane intersect with it? Justify your answer
1
Expert's answer
2020-12-01T06:31:11-0500

It is well-known that the equation of a two-sheet hyperboloid is


"-\frac{x^2}{a^2}-\frac{y^2}{b^2}+\frac{z^2}{c^2}=1"


Since the equation "1+2x^2+9y^2=3z^2" is equivalent to "-2x^2-9y^2+3z^2=1", we conclude that the concoid "1+2x^2+9y^2=3z^2" is a two-sheet hyperboloid.


If "z=0", then the equation "2x^2+9y^2=-1" has no real solutions, and therefore, the "XY"-plane does not intersect with the two-sheet hyperboloid 1+2x^2+9y^2=3z^2.